# FusionSAM: Latent Space driven Segment Anything Model for Multimodal Fusion and Segmentation

### v1

@@ Abstract
Figure 1: Comparative results of SAM2 and FusionSAM under different multimodal datasets: FusionSAM demonstrates superior boundary accuracy and structural completeness, while SAM2 struggles with misclassifications and unclear boundaries. To ensure fairness, the input is the fusion feature map of our method, annotations for points and boxes prompts are illustrated.
Introduction
@@
Accurate and comprehensive scene understanding is crucial for autonomous driving (Zhang and Demiris 2023). Due to the limitations of sensor imaging devices, no single modality sensor can independently provide a complete description of the scene (Zhou et al. 2024; Xue and Marculescu 2023; Cao et al. 2023b; Xu, Yuan, and Ma 2023). For instance, infrared sensors capture thermal radiation information, highlighting objects of interest such as pedestrians and vehicles (Bellagente et al. 2024). Conversely, visible light sensors capture reflected light, generating scenes rich in texture details (Liu et al. 2024). By combining these modalities, complementary details that might be missed by individual sensors can be captured, enhancing the model’s ability to perform semantic segmentation of the complete scene (Cao et al. 2023a). Therefore, the fusion of infrared and visible light images has become a mainstream solution for improving scene understanding and semantic segmentation. However, current semantic segmentation models struggle to comprehend densely packed elements in multimodal driving scenes, failing to fully represent the captured information for better subsequent segmentation results.
@@
In recent decades, advancements in semantic segmentation within deep learning have significantly propelled the understanding of multimodal scenes. Capturing efficient multimodal fusion representations is key to enhancing segmentation performance. A common approach involves feature-level fusion of infrared and visible light images using Convolutional Neural Networks (CNNs) to extract rich semantic representations, but the local constraints of CNNs make it challenging to effectively merge information from different modalities. As an alternative, Transformer architectures, with their attention mechanisms and ability to model long-range dependencies, facilitate better global fusion and utilization of complementary information (Li et al. 2023; Zhang et al. 2024). However, pure transformer architectures lack the flexibility required for scene understanding, especially in autonomous driving scenarios where elements are densely packed (Cao et al. 2023c), and edge textures of segmented categories are blurred due to varying lighting conditions and nighttime environments. Without intermediate fine-tuning guidance to focus on critical regions, segmentation distortions can occur, hindering better scene parsing. The Segment Anything Model (SAM) has emerged as a transformative method for single-modal natural scene segmentation due to its flexible prompting architecture (Ravi et al. 2024; Kirillov et al. 2023). Remarkably, the prompt architecture of SAM enhances the model’s ability to focus on detailed features. Through the guiding mechanism of prompts, SAM can more effectively direct the segmentation process compared to transformers that lack fine-tuned control. This is crucial for the dense element segmentation required in autonomous driving scenarios. However, SAM has not yet been extensively studied in the realm of multimodal fusion.
@@
To address these challenges, we innovatively propose FusionSAM, a Latent Space driven Segment Anything Model for Multi-Modal Fusion and Segmentation, which endows SAM with efficient multimodal image fusion and segmentation capabilities. Specifically, we first capture latent space feature embeddings of the two modalities through vector quantization to obtain efficient downsampled representations. Then, we establish long-range dependencies between the modalities using a cross-attention-based inter-domain fusion module, capturing comprehensive information as fusion features to guide precise pixel-level segmentation. To the best of our knowledge, this is the first study to apply the SAM to multimodal visual segmentation tasks in natural images, and it outperforms current state-of-the-art methods as shown in Figure 1. Our main contributions are as follows:
@@
    We propose a novel FusionSAM framework that includes the Latent Space Token Generation (LSTG) and Fusion Mask Prompting (FMP) Module. By capturing latent space representations through vector quantization and performing cross-domain fusion of these features, we generate precise segmentation prompts.
@@
    Extensive experiments on public datasets and benchmarks show that FusionSAM significantly outperforms state-of-the-art methods, including SAM and SAM2, in multimodal autonomous driving scenarios, achieving a notable 3.9% improvement in segmentation IoU, validating its effectiveness and robustness.
@@
Figure 2: Overview of FusionSAM framework for multimodal visual segmentation, which enhances multimodal visual understanding on the original SAM architecture. The main improvements include Latent Space Token Generation (LSTG) Module and Fusion Mask Prompting (FMP) Module. All but the Image encoder of architecture participate in the training phase.
Related Work
Segment Anything Model (SAM)
@@
The SAM enables efficient object segmentation through simple prompt embeddings, like points or bounding boxes, guiding the model to focus on specific regions (Ke et al. 2023; Schön et al. 2024; Ren et al. 2024; Shen, Li, and Wang 2024; Wang et al. 2023; Ma et al. 2024; Huang et al. 2024). Derived methods in single-modality segmentation include RobustSAM (Chen et al. 2024) by Chen et al., which improved SAM’s performance on low-quality images, and Crowd-SAM (Cai et al. 2024) by Cai et al., which enhanced segmentation in crowded scenes with an Efficient Prompt Sampler and Part-Whole Discriminator Network. SAM has also been adapted for cross-modal tasks in fields like medical imaging and remote sensing. For example, Pandey et al. used YOLOv8 and SAM for cross-modal segmentation (Pandey, Chen, and Dam 2023), while Yan et al. introduced RingMo-SAM (Yan et al. 2023) for segmenting optical and SAR data. However, these methods only linearly adapt SAM for multimodal tasks, missing the full potential of multimodal features. They also overlook SAM’s powerful prompting architecture, which could better activate multimodal fusion features during training to guide segmentation. Our proposed FusionSAM, on the other hand, captures latent space representations through vector quantization, enabling comprehensive cross-domain fusion and using these features as precise segmentation prompts.
Multi-Modality Image Fusion
@@
In autonomous driving, integrating various sensors is essential for accurate scene understanding, as single-modality data is insufficient (Liang et al. 2022; Li et al. 2024; Zhang et al. 2021; Sun et al. 2022; Wang et al. 2020a, 2022; Zhang, Ye, and Xiao 2020). Wang et al. proposed AsymFusion (Wang et al. 2020b), which enhances multimodal feature interaction using a dual-branch structure with asymmetric fusion blocks. Zhang et al. developed MRFS (Zhang et al. 2024), combining CNN-based Interactive Gated Mixed Attention with transformer-based Progressive Cycle Attention to overcome bottlenecks in infrared-visible fusion. Feng et al. introduced MAF-Net (Feng et al. 2022), which effectively segments road potholes by fusing RGB and disparity data. Ma et al. proposed SwinFusion (Ma et al. 2022), leveraging cross-domain long-range learning and Swin Transformer for global information integration and complementary feature extraction. Most existing methods rely on convolutional networks or transformers, which struggle with global information extraction and flexible segmentation in dense scenes. To overcome these limitations, we apply multimodal fusion within SAM, using its flexible prompting to enhance segmentation in complex autonomous driving scenarios.
Proposed Method
Problem Formulation
@@
For the task of multimodal image fusion, we first assume visible image $I_{1}\in\mathbb{R}^{H\times W\times C_{in}}$ and infrared image $I_{2}\in\mathbb{R}^{H\times W\times C_{in}}$, where the two source images from different domains are aligned. Let $H$, $W$, and $C_{in}$ denote the height, width, and channel number of input images, respectively. To achieve pixel-level segmentation, we design an interactive neural network for fusion and segmentation, and optimize the model to find a set of optimal parameters. The optimization model is formulated as follows:
$\min_{\boldsymbol{\omega}_{\mathrm{f}},\boldsymbol{\omega}_{\mathrm{s}}}f_{\mathrm{f}}\left(I_{f},\Phi\left(I_{1},I_{2};\boldsymbol{\omega}_{\mathrm{f}}\right)\right)+f_{\mathrm{s}}\left(I_{s},\Psi\left(I_{1},I_{2};\boldsymbol{\omega}_{\mathrm{s}}\right)\right),$ (1)
@@
$I_{f}\in\mathbb{R}^{H\times W\times C_{\mathrm{in}}}$ and $I_{s}\in\mathbb{R}^{H\times W\times C_{\mathrm{in}}}$ represent the fusion map and segmentation result, produced by the fusion network $\Phi$ and segmentation network $\Psi$ with learnable parameters $\boldsymbol{\omega}_{\mathrm{f}}$ and $\boldsymbol{\omega}_{\mathrm{s}}$. The functions $f_{\mathrm{f}}(\cdot)$ and $f_{\mathrm{s}}(\cdot)$ correspond to the objective functions for fusion and segmentation, measuring the discrepancies between the predictions and their respective targets. As shown in the upper right corner of Figure 2, the optimization goal is to minimize these functions, avoid falling into the local optimal of a single task, and ensure effective fusion and accurate segmentation to achieve global optimal.
Fusion Segment Anything Model
@@
We propose FusionSAM, which promotes image fusion while retaining the segmentation capability of SAM architecture. We use the adapter strategy to give the original SAM architecture the ability of hidden space representation embedding and cross-modal consistency fusion so that its segmentation performance has been greatly improved.
Model Overview.
@@
Figure 2 presents an overview of the proposed FusionSAM. The key contribution of FusionSAM is its Latent Space Token Generation (LSTG) and Fusion Mask Prompting (FMP) modules. Unlike methods that fine-tune or add adapters to SAM and SAM2, FusionSAM’s strength lies in its rigorous and well-considered approach to efficient multimodal fusion and segmentation.
Training.
@@
To train FusionSAM, we first generate efficient fused modality representations, which are then input into the model. Initially, a vector encoder creates latent space representations for both modalities, followed by cross-attention-guided fusion to achieve a comprehensive representation. Unlike the original SAM, we modify the input tokens for segmentation into Full-fledged Output Tokens (FOT). These FOTs, along with the prompt token, are processed through the SAM decoding layers to generate the segmentation mask.
@@
Inference.
In the FusionSAM framework, the ViT-driven image encoder is not involved in training, it is solely used for inference to generate inputs for the mask decoder.
Latent Space Token Generation.
In our multimodal image fusion and segmentation approach, the LSTG module effectively transforms complex input data from visible and infrared modalities into structured latent space representations. This transformation is essential for the efficient integration of diverse information sources. By drawing inspiration from Vector Quantized Generative Adversarial Networks, we enhance our model’s capability to capture and fuse complementary features from both modalities, thereby improving the performance of multimodal tasks.
@@
$z_{i}=\mathcal{E}(I_{i})\in\mathbb{R}^{h\times w\times d_{c}}.$ (2)
@@
The quantization process translates the encoder outputs $z_{i}$ into discrete representations using a learned codebook $\mathcal{C}$, aligning and structuring diverse features from both modalities for effective fusion:
$I_{i}^{q}=\operatorname{Quant}(z_{i})=\left(\operatorname{argmin}_{c_{k}\in\mathcal{C}}\|z_{ij}-c_{k}\|\right)\in\mathbb{R}^{h\times w\times d_{c}}.$ (3)
@@
The decoder $\mathcal{G}$ reconstructs the original images from these quantized representations, ensuring that the fused representation retains the high fidelity and rich detail necessary for accurate segmentation:
$\hat{I}_{i}=\mathcal{G}(I_{i}^{q})=\mathcal{G}(\operatorname{Quant}(\mathcal{E}(I_{i}))).$ (4)
@@
$\mathcal{L}_{\text{rec}}=\sum_{i}\|I_{i}-\hat{I}_{i}\|^{2},$ (5)
$\mathcal{L}_{\text{commit}}=\sum_{i}\|\operatorname{sg}[z_{i}]-I_{i}^{q}\|^{2}_{2}+\beta\|\operatorname{sg}[I_{i}^{q}]-z_{i}\|^{2}_{2},$ (6)
@@
$\mathcal{L}_{\text{perc}}=\sum_{i}\|\Phi(I_{i})-\Phi(\hat{I}_{i})\|^{2}.$ (7)
$\mathcal{L}_{\text{adv}}=\sum_{i}\left(\log\mathcal{D}(I_{i})+\log(1-\mathcal{D}(\hat{I}_{i}))\right),$ (8)
$\mathcal{D}$ is the discriminator network used in the adversarial learning framework, distinguishing between real and generated data. These enhancements ensure that the model captures both low-level detail and high-level semantic information, which is crucial for effective multimodal segmentation. The overall optimization objective combines these elements:
$\min_{\mathcal{E},\mathcal{G},\mathcal{C}}\max_{\mathcal{D}}\sum_{i}\left[\mathcal{L}_{\text{rec}}+\alpha\mathcal{L}_{\text{perc}}+\beta\mathcal{L}_{\text{adv}}+\gamma\mathcal{L}_{\text{commit}}\right],$ (9)
where $\alpha$, $\beta$, and $\gamma$ are weighting factors that balance the contributions of each loss component, enhancing the model’s ability to perform well on multimodal tasks.
Figure 3: Qualitative demonstrations of different approaches on the MFNet dataset.
@@
Ours 98.8 89.8 74.0 75.8 69.6 60.2 0.0 37.8 61.4 63.0
Table 1: Results of quantitative segmentation on the test set of MFNet dataset.
Fusion Mask Prompting Module.
The FMP module is designed to effectively synthesize latent space representations from visible and infrared modalities, enabling comprehensive scene understanding in autonomous driving scenarios. This module integrates information from different domains presented by each modality into a unified fusion mask. By leveraging the rich and comprehensive features present in the fusion representation as prompts, FMP module provides flexible fine-tuning guidance for the segmentation process, leading to improved segmentation performance. For instance, if multimodal fusion feature map contains complete information, using local area features as point prompts during training can further enhance the model’s segmentation accuracy.
Specifically, the FMP module begins with a cross-domain fusion unit that employs cross-attention mechanisms to establish long-range dependencies between different modality domains. This facilitates the exchange of Queries (Q), Keys (K) , and Values (V) across domains, ensuring the complete fusion of multimodal features. This process ensures that the fusion mask captures comprehensive interactions between the latent representations $I_{1}^{q}$ and $I_{2}^{q}$, enhancing the segmentation process by focusing on critical, contextually relevant features that are essential for understanding dense and complex scenes. The inter-domain mechanism is defined as follows:
$\left\{Q_{1},K_{1},V_{1}\right\}=\left\{I_{1}^{q}W_{Q1},I_{1}^{q}W_{K1},I_{1}^{q}W_{V1}\right\},$ (10)
$\left\{Q_{2},K_{2},V_{2}\right\}=\left\{I_{2}^{q}W_{Q2},I_{2}^{q}W_{K2},I_{2}^{q}W_{V2}\right\},$
$z_{1}^{\prime}=\text{LN}\left(\text{softmax}\left(\frac{Q_{1}K_{2}^{T}}{\sqrt{d_{k}}}\right)V_{2}\right)+Q_{1},$ (11)
$z_{2}^{\prime}=\text{LN}\left(\text{softmax}\left(\frac{Q_{2}K_{1}^{T}}{\sqrt{d_{k}}}\right)V_{1}\right)+Q_{2},$
the LN($\cdot$) is the layer normalization, whcih always performed after feed forward network, the outputs $z_{1}^{\prime}$ and $z_{2}^{\prime}$ represent the globally fused features, which are then processed through a convolutional layer, generating a fused representation $z_{c}$ that encapsulates the essential information from both modalities. This fused representation serves as the initial fusion mask, guiding the segmentation by highlighting the regions of interest identified through the cross-domain fusion process.

### v2

# FusionSAM: Visual Multi-Modal Learning with Segment Anything Model

@@ Abstract.
Scene understanding, Computer vision
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736973 isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Scene understanding
@@
Figure 1. Comparative results of SAM2 and FusionSAM under the MFNet and FMB datasets: FusionSAM demonstrates superior boundary accuracy and structural completeness, while SAM2 struggles with misclassifications and unclear boundaries. To ensure fairness, the input is the fusion feature map of our method, annotations for points and boxes prompts are shown.
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15523136.
1. Introduction
@@
Accurate and comprehensive scene understanding is crucial for scene understanding and autonomous driving (Zhang and Demiris, 2023; Min et al., 2024; Schön et al., 2021; Liao et al., 2024; Jiang et al., 2023). Due to the limitations of sensor imaging devices, no single modality sensor can independently provide a complete description of the scene (Zhou et al., 2024; Xue and Marculescu, 2023; Cao et al., 2023c; Xu et al., 2023). For instance, infrared sensors capture thermal radiation information, highlighting objects of interest such as pedestrians and vehicles (Bellagente et al., 2024; Sheinin et al., 2024; Erlenbusch et al., 2023; Jiang and Chen, 2024; Zhang et al., 2025; Zhao et al., 2024; Yi et al., 2024). Conversely, visible light sensors capture reflected light, generating scenes rich in texture details (Liu et al., 2024; Pendota and Channappayya, 2024). By combining these modalities, complementary details that might be missed by individual sensors can be captured, enhancing the model’s ability to perform semantic segmentation of the complete scene (Cao et al., 2023a; Mayr et al., 2024; Li et al., 2024a; Zheng et al., 2024; Ma et al., 2022). Therefore, the fusion of infrared and visible light images has become a mainstream solution for improving scene understanding and semantic segmentation. However, current semantic segmentation models struggle to comprehend densely packed elements in multimodal driving and other modality fusion scenes, failing to fully represent the captured information for better subsequent segmentation results.
@@
In recent decades, advancements in semantic segmentation within deep learning have significantly propelled the understanding of multimodal scenes. Capturing efficient multimodal fusion representations is key to enhancing segmentation performance. A common approach involves feature-level fusion of infrared and visible light images using Convolutional Neural Networks (CNNs) to extract rich semantic representations, but the local constraints of CNNs make it challenging to effectively merge information from different modalities. As an alternative, Transformer architectures, with their attention mechanisms and ability to model long-range dependencies, facilitate better global fusion and utilization of complementary information(Li et al., 2023b; Zhang et al., 2024b; Peng et al., 2024; Rahman et al., 2024; Yang et al., 2021).
However, pure transformer architectures lack the flexibility required for scene understanding, especially in autonomous driving scenarios where elements are densely packed (Cao et al., 2023b; Huang et al., 2023; Ye et al., 2022), and edge textures of segmented categories are blurred due to varying lighting conditions and nighttime environments. Without intermediate fine-tuning guidance to focus on critical regions, segmentation distortions can occur, hindering better scene parsing. The Segment Anything Model (SAM) has emerged as a transformative method for single-modal natural scene segmentation due to its flexible prompting architecture (Ravi et al., 2024; Kirillov et al., 2023; Wang et al., 2024b). Remarkably, the prompt architecture of SAM enhances the model’s ability to focus on detailed features. Through the guiding mechanism of prompts, SAM can more effectively direct the segmentation process compared to transformers that lack fine-tuned control. This is crucial for the dense element segmentation required in autonomous driving scenarios. However, SAM has not yet been extensively studied in the realm of multimodal fusion.
@@
To address the challenges of traditional approaches that directly input linearly fused results into SAM, which often leads to excessive redundancy in the fused modalities and weakens the effective information from each modality, an important issue arises: how to efficiently fuse the modalities while ensuring the fusion features are as comprehensive as possible, and use these features as effective segmentation prompts for SAM. To overcome this, we innovatively propose FusionSAM, a Latent Space driven Segment Anything Model for Multi-Modal Fusion and Segmentation, which endows SAM with efficient multimodal image fusion and segmentation capabilities. Specifically, we first capture latent space feature embeddings of the two modalities through vector quantization to obtain efficient downsampled representations. Then, we establish long-range dependencies between the modalities using a cross-attention-based inter-domain fusion module, capturing comprehensive information as fusion features to guide precise pixel-level segmentation. To the best of our knowledge, this is the first study to apply SAM to multimodal visual segmentation tasks in natural images, and it outperforms current state-of-the-art methods, as shown in Figure 1. Our main contributions are as follows:
@@
    We propose a novel FusionSAM framework that includes the Latent Space Token Generation (LSTG) and Fusion Mask Prompting (FMP) Module. By capturing latent space representations through vector quantization and performing cross-domain fusion of these features, we generate precise segmentation.
@@
    Extensive experiments on public datasets and benchmarks show that FusionSAM significantly outperforms state-of-the-art methods, including SAM and SAM2, in multimodal autonomous driving multi-modal visual scenes (RGB, Infrared, Depth) scenarios, achieving a notable 4.1% improvement in segmentation IoU, validating its effectiveness and robustness. The performance is also optimized in other multi-modal visual scenes.
@@
Figure 2. Overview of FusionSAM framework for multimodal visual segmentation, which enhances multimodal visual understanding on the original SAM architecture. The main improvements include Latent Space Token Generation (LSTG) Module and Fusion Mask Prompting (FMP) Module. Unlike the traditional SAM architecture, the Fusion Point Prompt Encoder in FusionSAM extracts point prompts directly from the fusion features, enabling more comprehensive and precise guidance for segmentation. All parts of the architecture except the image encoder participate in the training phase.
2. Related Work
2.1. Segment Anything Model (SAM)
@@
The SAM enables efficient object segmentation through simple prompt embeddings, like points or bounding boxes, guiding the model to focus on specific regions (Ke et al., 2023; Schön et al., 2024; Ren et al., 2024a; Shen et al., 2024a; Wang et al., 2023b; Ma et al., 2024; Huang et al., 2024; Wei et al., 2024; Ren et al., 2024b; Li et al., 2023c). Derived methods in single-modality segmentation include RobustSAM (Chen et al., 2024) by Chen et al., which improved SAM’s performance on low-quality images, and Crowd-SAM (Cai et al., 2024) by Cai et al., which enhanced segmentation in crowded scenes with an Efficient Prompt Sampler and Part-Whole Discriminator Network. SAM has also been adapted for cross-modal tasks in fields like medical imaging and remote sensing. For example, Pandey et al. used YOLOv8 and SAM for cross-modal segmentation (Pandey et al., 2023), while Yan et al. introduced RingMo-SAM (Yan et al., 2023) for segmenting optical and SAR data. However, these methods only linearly adapt SAM for multimodal tasks, missing the full potential of multimodal features. They also overlook SAM’s powerful prompting architecture, which could better activate multimodal fusion features during training to guide segmentation. Our proposed FusionSAM, captures latent space representations through vector quantization, enabling comprehensive cross-domain fusion and using these features as precise segmentation prompts.
2.2. Multi-Modality Image Fusion
@@
In autonomous driving, integrating various sensors is essential for accurate scene understanding, as single-modality data is insufficient (Liang et al., 2022; Li et al., 2024b; Zhang et al., 2021; Sun et al., 2022; Wang et al., 2020a, 2022; Zhang et al., 2020, 2024a; Wang et al., 2024a, 2023a; Ando et al., 2023; Li et al., 2023a). Wang et al. proposed AsymFusion (Wang et al., 2020b), which enhances multimodal feature interaction using a dual-branch structure with asymmetric fusion blocks. Zhang et al. developed MRFS (Zhang et al., 2024b), combining CNN-based Interactive Gated Mixed Attention with transformer-based Progressive Cycle Attention to overcome bottlenecks in infrared-visible fusion. Feng et al. introduced MAF-Net (Feng et al., 2022), which effectively segments road potholes by fusing RGB and disparity data. Ma et al. proposed SwinFusion (Ma et al., 2022), leveraging cross-domain long-range learning and Swin Transformer for global information integration and complementary feature extraction. Most existing methods rely on convolutional networks or transformers, which struggle with global information extraction and flexible segmentation in dense scenes. To overcome these limitations, we apply multimodal fusion within SAM, using its flexible prompting to enhance segmentation in complex autonomous driving scenarios.
3. Methodology
3.1. Problem Formulation
@@
For the task of multimodal image fusion, we first assume visible image $I_{1}\in\mathbb{R}^{H\times W\times C_{in}}$ and infrared image $I_{2}\in\mathbb{R}^{H\times W\times C_{in}}$, where the two source images from different domains are aligned. Let $H$, $W$, and $C_{in}$ denote the height, width, and channel number of input images, respectively. As shown in the upper right corner of Figure 2, to achieve pixel-level segmentation, we design an interactive neural network for fusion and segmentation, and optimize the model to find a set of optimal parameters. The optimization model is formulated as follows:
(1) $\min_{\boldsymbol{\omega}_{\mathrm{f}},\boldsymbol{\omega}_{\mathrm{s}}}f_{\mathrm{f}}\Big{(}I_{f},\Phi\big{(}I_{1},I_{2};\boldsymbol{\omega}_{\mathrm{f}}\big{)}\Big{)}+f_{\mathrm{s}}\Big{(}I_{s},\Psi\big{(}I_{1},I_{2};\boldsymbol{\omega}_{\mathrm{s}}\big{)}\Big{)},$
@@
$I_{f}\in\mathbb{R}^{H\times W\times C_{\mathrm{in}}}$ and $I_{s}\in\mathbb{R}^{H\times W\times C_{\mathrm{in}}}$ represent the fusion map and segmentation result, produced by the fusion network $\Phi$ and segmentation network $\Psi$ with learnable parameters $\boldsymbol{\omega}_{\mathrm{f}}$ and $\boldsymbol{\omega}_{\mathrm{s}}$. The functions $f_{\mathrm{f}}(\cdot)$ and $f_{\mathrm{s}}(\cdot)$ correspond to the objective functions for fusion and segmentation, measuring the discrepancies between the predictions and their respective targets.
3.2. Fusion Segment Anything Model
@@
We propose FusionSAM, which enhances image fusion while preserving the segmentation capability of the SAM architecture. By integrating a fusion module that enables latent space representation embedding and cross-modal consistency fusion into the original SAM architecture, so that its performance will be greatly improved.
3.2.1. Model Overview
@@
Figure 2 presents an overview of the proposed FusionSAM. The key contribution of FusionSAM is its Latent Space Token Generation (LSTG) and Fusion Mask Prompting (FMP) modules. Unlike methods that fine-tune or add adapters to SAM and SAM2, FusionSAM’s strength lies in its rigorous and well-considered approach to efficient multimodal fusion and segmentation. This efficiency is achieved by fusing compact and comprehensive latent space representations of both modalities, rather than the original large-scale images, enabling more thorough and effective fusion.
@@
Training. To train FusionSAM, we generate efficient fused modality representations, which are then input into model. Initially, a vector encoder creates latent space representations for both modalities, followed by cross-attention-guided fusion to achieve a comprehensive representation. Unlike the original SAM, we modify the input tokens for segmentation into Full-fledged Output Tokens (FOT), which are enhanced versions of the latent representations designed to capture the full spectrum of fused features for segmentation. These FOTs, along with the prompt token, are processed through the SAM decoding layers to generate the segmentation mask.
@@
Inference. In the FusionSAM framework, the ViT-driven image encoder is not involved in training, it is solely used for inference to generate inputs for the mask decoder.
3.2.2. Latent Space Token Generation
In our multimodal image fusion and segmentation approach, the LSTG module effectively transforms complex input data from visible and infrared modalities into structured latent space representations. This transformation is essential for the efficient integration of diverse information sources. By drawing inspiration from Vector Quantized Generative Adversarial Networks (VQGAN) (Esser et al., 2021), we enhance our model’s capability to capture and fuse complementary features from both modalities, thereby improving the performance of multimodal tasks.
@@ 3.2.2. Latent Space Token Generation
(2) $z_{i}=\mathcal{E}(I_{i})\in\mathbb{R}^{h\times w\times d_{c}}.$
@@
The quantization process translates encoder outputs $z_{i}$ into discrete representations using a learned codebook $\mathcal{C}$, aligning and structuring diverse features from both modalities for fusion:
(3) $I_{i}^{q}=\operatorname{Quant}(z_{i})=\left(\operatorname{argmin}_{c_{k}\in\mathcal{C}}\|z_{ij}-c_{k}\|\right)\in\mathbb{R}^{h\times w\times d_{c}}.$
@@
The decoder $\mathcal{G}$ reconstructs the original images from these quantized representations, ensuring that the fusion representation obtained in the subsequent fusion process retains the high fidelity and rich details required for accurate segmentation:
(4) $\hat{I}_{i}=\mathcal{G}(I_{i}^{q})=\mathcal{G}(\operatorname{Quant}(\mathcal{E}(I_{i}))).$
@@
(5) $\mathcal{L}_{\text{rec}}=\sum_{i}\|I_{i}-\hat{I}_{i}\|^{2},$
(6) $\mathcal{L}_{\text{commit}}=\sum_{i}\|\operatorname{sg}[z_{i}]-I_{i}^{q}\|^{2}_{2}+\beta\|\operatorname{sg}[I_{i}^{q}]-z_{i}\|^{2}_{2},$
@@
(7) $\mathcal{L}_{\text{perc}}=\sum_{i}\|\Phi(I_{i})-\Phi(\hat{I}_{i})\|^{2}.$
(8) $\mathcal{L}_{\text{adv}}=\sum_{i}\left(\log\mathcal{D}(I_{i})+\log(1-\mathcal{D}(\hat{I}_{i}))\right),$
$\Phi(\cdot)$ represents the feature map of the reconstructed image in VQGAN, $\mathcal{D}$ is the discriminator network used in the adversarial learning framework, distinguishing between real and generated data. These enhancements ensure that the model captures both low-level detail and high-level semantic information, which is crucial for effective multimodal segmentation.
As a result, we choose vector quantization in the Latent Space Token Generation (LSTG) module due to its ability to generate compact and semantically aligned latent representations across modalities. Unlike raw features, VQ discretizes inputs into a codebook, making the fused features more structured and suitable for prompt extraction in the SAM architecture. We found that this approach not only reduces spatial redundancy but also preserves modality-specific information in a consistent latent space, enabling more efficient and robust fusion.
Figure 3. A visual demonstration of different methods on the MFNet dataset.
Table 1. Results of quantitative segmentation on the test set of MFNet dataset.
@@
Ours 98.7 89.6 76.2 75.8 69.6 50.1 10.2 47.6 61.4 64.4
3.2.3. Fusion Mask Prompting Module
As shown in Algorithnm 1, this module integrates information from different domains presented by each modality into a unified fusion mask based on the SwinFusion. By leveraging the rich and comprehensive features present in the fusion representation as prompts, FMP module provides flexible fine-tuning guidance for the segmentation process, leading to improved segmentation performance. For instance, if multimodal fusion feature map contains complete information, using local area features as point prompts during training can further enhance the model’s segmentation accuracy.
Algorithm 1 Fusion Mask Prompting (FMP) Module Fusion.
1:RGB feature map $\mathbf{F}_{\text{RGB}}$, Thermal feature map $\mathbf{F}_{\text{T}}$
2:Fused feature $\mathbf{F}_{\text{out}}$
3:Step 1: Self-Attention on Each Modality
4:$\mathbf{F}_{\text{RGB}}^{\text{self}}\leftarrow\text{SelfAttention}(\mathbf{F}_{\text{RGB}})$
5:$\mathbf{F}_{\text{T}}^{\text{self}}\leftarrow\text{SelfAttention}(\mathbf{F}_{\text{T}})$
6:Step 2: Cross-Attention with Query from RGB and Key / Value from Thermal
7:$\mathbf{Q}\leftarrow\mathbf{F}_{\text{RGB}}^{\text{self}}\cdot\mathbf{W}_{Q}$
8:$\mathbf{K}\leftarrow\mathbf{F}_{\text{T}}^{\text{self}}\cdot\mathbf{W}_{K}$
9:$\mathbf{V}\leftarrow\mathbf{F}_{\text{T}}^{\text{self}}\cdot\mathbf{W}_{V}$
10:$\mathbf{A}\leftarrow\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)$
11:$\mathbf{F}_{\text{fused}}\leftarrow\mathbf{A}\cdot\mathbf{V}$
12:Step 3: Residual Connection and Layer Normalization
13:$\mathbf{F}_{\text{out}}\leftarrow\text{LayerNorm}(\mathbf{F}_{\text{fused}}+\mathbf{F}_{\text{RGB}}^{\text{self}})$ return $\mathbf{F}_{\text{out}}$
Specifically, FMP module begins with a cross-domain fusion unit that employs cross-attention mechanisms to establish long-range dependencies between different modality domains. This facilitates the exchange of Queries ($Q$), Keys ($K$) , and Values ($V$) across domains, ensuring the complete fusion of multimodal features. This process ensures that the fusion mask captures comprehensive interactions between the latent representations $I_{1}^{q}$ and $I_{2}^{q}$, enhancing the segmentation process by focusing on critical, contextually relevant features that are essential for understanding dense and complex scenes. The inter-domain mechanism is defined as:
(9) $\left\{Q_{1},K_{1},V_{1}\right\}=\left\{I_{1}^{q}W_{Q1},I_{1}^{q}W_{K1},I_{1}^{q}W_{V1}\right\},$
$\left\{Q_{2},K_{2},V_{2}\right\}=\left\{I_{2}^{q}W_{Q2},I_{2}^{q}W_{K2},I_{2}^{q}W_{V2}\right\}.$
(10) $z_{1}^{\prime}=\text{LN}\left(\text{softmax}\left(\frac{Q_{1}K_{2}^{T}}{\sqrt{d_{k}}}\right)V_{2}\right)+Q_{1},$
$z_{2}^{\prime}=\text{LN}\left(\text{softmax}\left(\frac{Q_{2}K_{1}^{T}}{\sqrt{d_{k}}}\right)V_{1}\right)+Q_{2}.$
Equations 9 and 10 present the process of the Fusion Conv of Figure 2. LN($\cdot$) is the layer normalization, whcih always performed after feed forward network, the outputs $z_{1}^{\prime}$ and $z_{2}^{\prime}$ represent the globally fused features, which are then processed through a convolutional layer, generating a fused representation $z_{c}$ that encapsulates the essential information from both modalities. This fused representation serves as the initial fusion mask, guiding the segmentation by highlighting the regions of interest identified through the cross-domain fusion process.